OpenStudy (anonymous):

how to write 47/40 as a decimal

OpenStudy (solomonzelman):

use your calculator to divide 47 by 40,\[1.175\]

OpenStudy (unklerhaukus):

47/40 =(40 + 7)/40 =1 + 7/40 =1 +(5+2)/40 =1 +5/40+2/40 =

OpenStudy (solomonzelman):

Well, the question asks to write it as a decimal, (you probably knew I'm going to say this)... It's a good way to brake it up though.

OpenStudy (unklerhaukus):

5/40, and 2/40 have cancelling factors so are is easier to convert to decimal than 7/40

OpenStudy (solomonzelman):

but you would have to add them, and this would give you the exact same result.

OpenStudy (unklerhaukus):

1 + 5/40 + 2/40 =1 + 1/8 + 1/20 =1 + 0.125 + 0.05 =1.175

OpenStudy (alekos):

i like that. doesnt require a calculator!

OpenStudy (solomonzelman):

well you can do it as 47/40 without a calculator too, but yes, that's an easier way to brake it up, if calculator is not allowed.